# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v2

@@ A.1.6 Summary of RLTO parameter values for figures
[…]
LABEL:Mfig:RLTOmodelF $10^{-5}$ $10$ $10^{2}$ Range Range 0
LABEL:Mfig:overepA Range NA Range Local optimum Local optimum 0
LABEL:Mfig:overepB Range NA Range Local optimum Local optimum 0
LABEL:Mfig:proteomA Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
LABEL:Mfig:proteomB Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
LABEL:Mfig:proteomC Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
S1A $10^{-5}$ $10$ $10^{2}$ Range Range 0
S1B $10^{-5}$ $10$ $10^{2}$ Range Range 0
S1C $10^{-5}$ $10$ $10^{2}$ Range Range 0
S1D $10^{-5}$ $10$ $10^{2}$ Range Range 0
@@
S5A Range NA Range Local optimum Local optimum 0
S5B $10^{-5}$ NA Range Local optimum 30 0.1
@@
S7A Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
S7B Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
S7C Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
S7D Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
S8 Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
S10A Range NA Range Local optimum Local optimum 0
S14B Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
Table S1: RLTO model parameters by figure. Range appears if a range of parameters is used. NA appears if the parameter value is irrelevant. Local optimum appears if the parameter values in optimized to maximize the growth rate. All calculations are independent of $\delta\tau$ and $L_{0}$.
Figure S2: Strategies for robustness to noise. We schematically represent the stochastic protein number as waves. To achieve robust function, protein levels must be above the threshold. Robustness can be achieved by two distinct strategies: (i) increased expression level and (ii) decreased noise.
A.2 Discussion: The fitness landscape of a trade-off.
The fitness landscape predicted by the RLTO model for representative parameters is shown in Fig. S1. The figure displays a number of important model phenomena: There is no growth for mean protein number $\mu_{p}$ below the threshold number $n_{p}$, and for high noise, $\mu_{p}$ must be in significant excess of $n_{p}$. Rapid growth can be achieved by the two mechanisms described in Fig. S2: (i) high expression levels ($\mu_{p}$) are required for high noise amplitude ( $\rm{CV}_{p}^{2}$ ) or (ii) lower expression levels coupled with lower noise. This trade-off leads to a ridge-like feature of nearly optimal models. The optimal fitness corresponds to a balance between increasing the mean protein number ($\mu_{p}$) and decreasing noise ( ${\rm CV}_{p}^{2}$ ). This optimal central dogma program strategy leads to significant overabundance.
@@
Figure S4: All these calculations are performed for protein cost $\lambda=100$ in order to give real numbers in molecules per cell. Panel A: Protein number versus protein threshold. At high expression levels, the protein number tracks the protein threshold; however, the one message rule forces the protein number to threshold for low expression levels. These calculations are $\lambda$ dependent. Panel B: Message number versus message threshold. At high expression levels, the message number tracks the message threshold; however, the one message rule forces the message number to a threshold close to $\mu_{m}=1$ for low expression levels. These calculations are $\lambda$ independent. Panel C & D: Message number versus translation efficiency. The optimal translation efficiency grows almost linearly with the optimal message number. The scaled translation efficiency ($\hat{\varepsilon}/\lambda$) is independent of $\lambda$ while the translation efficiency ($\hat{\varepsilon}$) is dependent on $\lambda$. The ratio $\varepsilon/\lambda$ has a second interpretation: the load ratio $R$. $R$ is defined as the metabolic cost of translation over transcription of the gene.
A.3 Methods: Central dogma optimization
A.3.1 Optimization of transcription and translation (eukaryotes)
@@
$\ln\frac{k}{k_{0}}=\textstyle-(\Lambda+E)\mu_{m}-\frac{1}{\ln 2}\gamma(\frac{\mu_{m}}{\ln 2},\frac{\phi_{p}}{E\ln 2}),$ (18)
where $\gamma$ is the regularized incomplete gamma function, which is the CDF of the gamma distribution and represents the probability of arrest due to gene $i$. We set the partial derivative of the growth rate with respect to message number equal to zero:
@@
in both Eqs. 19 and 20. Eq. 20 can now be solved for the optimal translation efficiency:
@@
Figure S5: Panel A: Overabundance as a function of message number. In the limit of high message number ($\mu_{m}$), the noise is small, and the optimal expression level is close to the threshold $\phi_{p}$, leading to minimal overabundance ($o\rightarrow 1$); however, as the message number approaches unity, the noise is comparable to the mean, driving vast overabundance ($o\gg 1$). Overabundance increases as the relative metabolic load decreases, and is essentially identical for constant (dashed) versus optimized translation efficiency (solid). To model overabundance in bacterial cells, we included the contribution of the noise floor with fixed translation efficiency (dotted curves), which increases overabundance, especially for high expression genes. Panel B: Optimal expression levels are buffered. The predicted fitness loss as a function of protein depletion level and message number for bacterial cells (including the noise floor). Due to the overabundance phenomenon, all proteins are buffered against depletion, but low-expression genes are particularly robust due to higher overabundance. The solid red line represents $1/o$, and predicts the range of depletion values for which cell growth is predicted. The dotted red line represents a three-fold depletion.
A.3.2 Optimization of message number only
@@
We compare this modified expression to the original for optimum overabundance as a function of message number in Fig. S5A and demonstrate that the two make nearly identical predictions.
A.3.3 Inclusion of the noise floor
@@
where $C_{0}=0.1$ for bacterial cells [4]. In spite of the addition of noise from the noise floor, the observed distribution of protein number is still well described by the gamma distribution [4]; however, we need to modify the statistical parameters to account for the noise floor. (See the definition of the statistical noise model in Sec. D.1.2.) The modified gamma parameters are:
@@
The qualitative effect of the noise floor is to increase the noise, especially for low-copy messages. Above $\mu_{m}=7$ messages, the noise is dominated by the noise floor. Increases in transcription above this point have little effect on reducing the noise. As a consequence, the overabundance stays high, even for high copy messages. We compare this modified expression to the original for optimum overabundance as a function of message number in Fig. S5A and demonstrate that bacterial cells are predicted to have much higher overabundance at high expression levels.
A.4 Discussion: Understanding the rationale for overabundance
@@
Should this strategy surprise us? No. We routinely use this strategy in our own lives. For instance, when baking an apple pie, we keep cheap ingredients, needed in small quantities, in vast overabundance (e.g. salt); whereas, the more expensive ingredients, needed in higher quantities (e.g. apples), are bought only as needed. (See Fig. LABEL:Mfig:principlesA.)
A.5 Discussion: RLTO predicts larger overabundance in bacteria.
There are two distinctive features of bacterial cells that could affect the model predictions: (i) the translation efficiency is constant [5] and bacterial gene expression is subject to a large-magnitude noise floor that increases the noise for high-expression genes [4]. The optimization of message number at fixed translation efficiency does result in a slightly modified optimization condition for the message number (Sec. A.3.2); however, the predicted overabundance is only subtly perturbed (Fig. S5A). In contrast, the noise floor increases the predicted overabundance, especially for high-expression proteins. As a result, the RLTO model predicts that the vast majority of bacterial proteins are expressed in significant overabundance. (See Fig. S5A.)
A.6 Discussion: RLTO predicts proteins are buffered to depletion.
A principle motivation for our analysis is the observation that many protein levels appear to be buffered. To explore the prediction of the RLTO model for protein depletion, we first computed the optimal message numbers and translation efficiencies for a range of protein thresholds. To model the effect of protein depletion, we computed the change in growth rate as function of protein depletion (equivalent to a reduction of the translation efficiency relative to the optimum.) The growth rate is shown in Fig. S5B for the RLTO model with parameters representative of a bacterial cell. (See Sec. A.3.3.)
@@
A.7 Results: Detailed development of load balancing
A.7.1 Prediction of the optimal load ratio.
@@ A.7 Results: Detailed development of load balancing
A.7.2 Measurements of the load ratio
@@
A.8 Results: Translation efficiency is predicted to increase with transcription.
@@
where $\hat{R}$ depends weakly on the relative load $\Lambda$. The RLTO model predicts that optimal partitioning of amplification between transcription (gain $\mu_{m}$) and translation (gain $\varepsilon$) has two important qualitative features: (i) As the message cost ($\lambda$) rises, the optimal translation efficiency increases in proportion. (ii) The optimal translation efficiency is also approximately proportional to message number ($\hat{\varepsilon}\propto\mu_{m}$). (See Fig. S7A.) Therefore, the RLTO model predicts that low expression levels should be achieved with low levels of transcription and translation, whereas high expression genes are achieved with high levels of both. We call this relation between optimal transcription and translation load balancing.
A.9 Results: RLTO predicts that message number responds to message cost
@@
Figure S6: Panel A: Message number decreases with increased relative load $\Lambda$. The optimal message number responds to changes in the message cost. The RLTO model predicts an approximate power-law relation (linear on a log-log plot) between message numbers. Panel B: A power-law relation is observed. To test whether central dogma regulation would adapt dynamically as predicted, we analyzed the relation between the yeast transcriptome under reference conditions and phosphate depletion (perturbed), which increases the message cost [6]. (Data from Ref. [7].) As predicted by the RLTO model, a global change in regulation is observed, which generates a power-law relation with scaling exponent $\alpha=0.837\pm 0.01$. The observed exponent is smaller than one, as predicted by an increased relative load $\Lambda$.
A.10 Results: RLTO predicts the yeast global regulatory response
@@
A.11 Methods: Prediction of the proteome fraction
A.11.1 Parameter-free prediction of proteome fraction.
@@ A.11 Methods: Prediction of the proteome fraction
A.11.2 RLTO: proteome fraction
@@
which can be used to predict the proteome fraction:
$\hat{\Phi}_{pi}\equiv\textstyle\frac{\hat{\mu}_{pi}}{\sum_{j}\hat{\mu}_{pj}},$ (42)
@@
$C\equiv\textstyle\sum_{i}\left[\mu_{m}p_{\Gamma}(\mu_{m}|\frac{\mu_{m}}{\ln 2},o\ln 2)|_{o=\hat{o}(\mu_{m})}\right]_{i},$ (43)
@@
$\hat{\Phi}_{pi}=C^{-1}\left[\mu_{m}p_{\Gamma}(\mu_{m}|\frac{\mu_{m}}{\ln 2},o\ln 2)|_{o=\hat{o}(\mu_{m})}\right]_{i},$ (44)
which generates the predicted solid curves shown in Fig. S7BCD.
A.11.3 Constant-translation-efficiency model: proteome fraction
@@
$\Phi_{pi}^{\prime}={C^{\prime}}^{-1}\mu_{mi},$ (46)
which generates the predicted dotted curves shown in Fig. S7BCD.
A.11.4 Sources of experimental data for proteome fraction analysis
@@
Figure S7: Panel A: Load balancing: the optimal translation efficiency increases with message number and cost. The ratio of the optimal translation efficiency ($\hat{\varepsilon}$) to the message cost ($\lambda$) is roughly independent of the relative load ($\Lambda$); therefore, the optimal translation efficiency is expected to be proportional to the message cost. The translation efficiency is also roughly proportional to the optimal message number $\hat{\mu}_{m}$. The ratio $\varepsilon/\lambda$ has a second interpretation: the load ratio $R$. $R$ is defined as the metabolic cost of translation over transcription of the gene. Gene proteome fraction versus message number for three model systems. Two models for relation between gene proteome fraction and message number are compared with observations: The RLTO model makes a parameter-free prediction of the optimal relation (optimal, solid red line). A second competing model is constant translation efficiency (constant, dotted red line). Panel B: Yeast proteome fraction. The RLTO prediction (solid) is superior to the constant-translation-efficiency prediction (dashed). Panel C: Mammalian proteome fraction. The RLTO prediction (solid) is superior to the constant-translation-efficiency prediction (dashed). Panel D: E. coli proteome fraction. In contrast, the constant-translation-efficiency prediction (dashed) is superior to RLTO prediction (solid).
A.12 Results: Load balancing is observed in eukaryotic cells
A non-trivial prediction of the RLTO Model is that translation efficiency and message number should be roughly proportional. Qualitatively, this strategy allows expression levels to be increased while distributing the added metabolic load between transcription, which reduces noise, and translation, which does not affect the noise. We predict the optimal translation efficiency versus message number which matches the observations in eukaryotic cells (Fig. S7BC). However, in E. coli, the translation efficiency and message number are not strongly correlated (Fig. S7D). Why does this organism appear not to load balance? We demonstrate that the observed translation efficiency is consistent with the RLTO model, augmented by a ribosome-per-message limit. (See Sec. A.14.) Hausser et al. have proposed just such a limit, based on the ribosome footprint of mRNA molecules [10]. (See Sec. A.16.) Although this augmented model is consistent with central dogma regulation in E. coli, it is not a complete rationale. This proposed translation-rate limit could be circumvented by increasing the lifetime of E. coli messages which would increase the translation efficiency. Why the message lifetime is as short as observed will require a more detailed E. coli-specific analysis.
A.13 Discussion: Relation between load balancing and previous results

### v3

@@ A.2.8 Summary of RLTO parameter values for figures
and panel: (No units) (molecules-1) (molecules) (molecules/cell cycle) (No units) (No units)
2A $10^{-5}$ $10$ $10^{2}$ Range Range 0
2C Range NA Range Local optimum Local optimum 0
4A Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
4B Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
4C Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
4D Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
5 Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
S2A $10^{-5}$ $10$ $10^{2}$ Range Range 0
S2B $10^{-5}$ $10$ $10^{2}$ Range Range 0
S2C $10^{-5}$ $10$ $10^{2}$ Range Range 0
S2D $10^{-5}$ $10$ $10^{2}$ Range Range 0
@@ 5 Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
S5 $10^{-5}$ NA Range Local optimum 30 0.1
@@
S7 Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
S9A Range NA Range Local optimum Local optimum 0
S12B Weak dependence ($10^{-5}$) NA Range Local optimum Local optimum 0
Table S1: RLTO model parameters by figure. Range appears if a range of parameters is used. NA appears if the parameter value is irrelevant. Local optimum appears if the parameter values in optimized to maximize the growth rate.
A.3 Discussion: The fitness landscape of a trade-off.
The fitness landscape predicted by the RLTO model for representative parameters is shown in Fig. S2. The figure displays a number of important model phenomena: There is no growth for mean protein number $\mu_{p}$ below the threshold number $n_{p}$, and for high noise, $\mu_{p}$ must be in significant excess of $n_{p}$. Rapid growth can be achieved by the two mechanisms: (i) high expression levels ($\mu_{p}$) are required for high noise amplitude ( $\rm{CV}_{p}^{2}$ ) or (ii) lower expression levels coupled with lower noise. This trade-off leads to a ridge-like feature of nearly optimal models. The optimal fitness corresponds to a balance between increasing the mean protein number ($\mu_{p}$) and decreasing noise ( ${\rm CV}_{p}^{2}$ ). This optimal central dogma program strategy leads to significant overabundance.
@@
Figure S4: Four perspectives on load balancing. All these calculations are performed for protein cost $\lambda=100$ in order to give real numbers in molecules per cell. Panel A: Protein number versus protein threshold. At high expression levels, the protein number tracks the protein threshold; however, the one message rule forces the protein number to threshold for low expression levels. These calculations are $\lambda$ dependent. Panel B: Message number versus message threshold. At high expression levels, the message number tracks the message threshold; however, the one message rule forces the message number to a threshold close to $\mu_{m}=1$ for low expression levels. These calculations are $\lambda$ independent. Panel C & D: Message number versus translation efficiency. The optimal translation efficiency grows almost linearly with the optimal message number. The scaled translation efficiency ($\hat{\varepsilon}/\lambda$) is independent of $\lambda$ while the translation efficiency ($\hat{\varepsilon}$) is dependent on $\lambda$. The ratio $\varepsilon/\lambda$ has a second interpretation: the load ratio $R$. $R$ is defined as the metabolic cost of translation over transcription of the gene.
A.4 Methods: Central dogma optimization
A.4.1 Optimization of transcription and translation (eukaryotes)
@@
$\frac{\Delta k}{k_{0}}=\textstyle-(\Lambda+\frac{\varepsilon}{N_{0}})\mu_{m}-\frac{1}{\ln 2}\gamma(\frac{\mu_{m}}{\ln 2},\frac{n_{p}}{\varepsilon\ln 2}),$ (54)
where $\gamma$ is the regularized incomplete gamma function, which is the CDF of the gamma distribution and represents the probability of arrest due to gene $i$. (Note that this equation is identical to Eq. 53 but with the gene subscript $i$ implicit.) We set the partial derivative of the growth rate with respect to message number equal to zero:
@@
in both Eqs. 55 and 56. Eq. 56 can now be solved for the optimal translation efficiency:
@@
Figure S5: Optimal expression levels are buffered. The predicted fitness loss as a function of protein depletion level and message number for bacterial cells (including the noise floor). Due to the overabundance phenomenon, all proteins are buffered against depletion, but low-expression genes are particularly robust due to higher overabundance. The solid red line represents $1/o$, and predicts the range of depletion values for which cell growth is predicted. The dotted red line represents a three-fold depletion.
A.4.2 Optimization of message number only
@@
We compare this modified expression to the original for optimum overabundance as a function of message number in Fig. 2C and demonstrate that the two make nearly identical predictions.
A.4.3 Inclusion of the noise floor
@@
where $C_{0}=0.1$ for bacterial cells [9]. In spite of the addition of noise from the noise floor, the observed distribution of protein number is still well described by the gamma distribution [9]; however, we need to modify the statistical parameters to account for the noise floor. (See the definition of the statistical noise model in Sec. A.1.2.) The modified gamma parameters are:
@@
The qualitative effect of the noise floor is to increase the noise, especially for low-copy messages. Above $\mu_{m}=7$ messages, the noise is dominated by the noise floor. Increases in transcription above this point have little effect on reducing the noise. As a consequence, the overabundance stays high, even for high copy messages. We compare this modified expression to the original for optimum overabundance as a function of message number in Fig. 2C and demonstrate that bacterial cells are predicted to have much higher overabundance at high expression levels.
A.5 Discussion: Understanding the rationale for overabundance
@@
A.6 Discussion: RLTO predicts larger overabundance in bacteria.
There are two distinctive features of bacterial cells that could affect the model predictions: (i) the translation efficiency is constant [22] and bacterial gene expression is subject to a large-magnitude noise floor that increases the noise for high-expression genes [9]. The optimization of message number at fixed translation efficiency does result in a slightly modified optimization condition for the message number (Sec. A.4.2); however, the predicted overabundance is only subtly perturbed (Fig. 2C). In contrast, the noise floor increases the predicted overabundance, especially for high-expression proteins. As a result, the RLTO model predicts that the vast majority of bacterial proteins are expressed in significant overabundance. (See Fig. 2C.)
A.7 Discussion: RLTO predicts proteins are buffered to depletion.
A principle motivation for our analysis is the observation that many protein levels appear to be buffered. To explore the prediction of the RLTO model for protein depletion, we first computed the optimal message numbers and translation efficiencies for a range of protein thresholds. To model the effect of protein depletion, we computed the change in growth rate as function of protein depletion (equivalent to a reduction of the translation efficiency relative to the optimum.) The growth rate is shown in Fig. S5 for the RLTO model with parameters representative of a bacterial cell. (See Sec. A.4.3.)
@@
A.8 Results: Detailed development of load balancing
A.8.1 Prediction of the optimal load ratio.
@@ A.8 Results: Detailed development of load balancing
A.8.2 Measurements of the load ratio
@@
A.9 Results: Translation efficiency is predicted to increase with transcription.
@@
where $\hat{R}$ depends weakly on the relative load $\Lambda$. The RLTO model predicts that optimal partitioning of amplification between transcription (gain $\mu_{m}$) and translation (gain $\varepsilon$) has two important qualitative features: (i) As the message cost ($\lambda$) rises, the optimal translation efficiency increases in proportion. (ii) The optimal translation efficiency is also approximately proportional to message number ($\hat{\varepsilon}\propto\mu_{m}$). (See Fig. 4A.) Therefore, the RLTO model predicts that low expression levels should be achieved with low levels of transcription and translation, whereas high expression genes are achieved with high levels of both. We call this relation between optimal transcription and translation load balancing.
A.10 Results: RLTO predicts that message number responds to message cost
@@
Figure S6: The message cost affects transciption genome wide. Panel A: Message number decreases with increased relative load $\Lambda$. The optimal message number responds to changes in the message cost. The RLTO model predicts an approximate power-law relation (linear on a log-log plot) between message numbers. Panel B: A power-law relation is observed. To test whether central dogma regulation would adapt dynamically as predicted, we analyzed the relation between the yeast transcriptome under reference conditions and phosphate depletion (perturbed), which increases the message cost [41]. (Data from Ref. [42].) As predicted by the RLTO model, a global change in regulation is observed, which generates a power-law relation with scaling exponent $\alpha=0.837\pm 0.01$. The observed exponent is smaller than one, as predicted by an increased relative load $\Lambda$.
A.11 Results: RLTO predicts the yeast global regulatory response
@@
A.12 Methods: Prediction of the proteome fraction
A.12.1 Parameter-free prediction of proteome fraction.
@@ A.12 Methods: Prediction of the proteome fraction
A.12.2 RLTO: proteome fraction
@@
which can be used to predict the proteome fraction (where we have restored the explicit gene $i$ subscript):
$\hat{\Phi}_{p,i}\equiv\textstyle\frac{\hat{\mu}_{p,i}}{\sum_{j}\hat{\mu}_{p,j}},$ (78)
@@
$C\equiv\textstyle\sum_{i}\left[\mu_{m,i}p_{\Gamma}(\mu_{m}|\frac{\mu_{m,i}}{\ln 2},o\ln 2)|_{o=\hat{o}(\mu_{m,i})}\right],$ (79)
@@
$\hat{\Phi}_{p,i}=C^{-1}\left[\mu_{m,i}p_{\Gamma}(\mu_{m,i}|\frac{\mu_{m,i}}{\ln 2},o\ln 2)|_{o=\hat{o}(\mu_{m,i})}\right],$ (80)
which generates the predicted solid curves shown in Fig. 4BCD.
A.12.3 Constant-translation-efficiency model: proteome fraction
@@
$\Phi_{p,i}^{\prime}={C^{\prime}}^{-1}\mu_{m,i},$ (82)
which generates the predicted dotted curves shown in Fig. 4BCD.
A.12.4 Sources of experimental data for proteome fraction analysis
@@
A.13 Results: Load balancing is observed in eukaryotic cells
A non-trivial prediction of the RLTO Model is that translation efficiency and message number should be roughly proportional. Qualitatively, this strategy allows expression levels to be increased while distributing the added metabolic load between transcription, which reduces noise, and translation, which does not affect the noise. We predict the optimal translation efficiency versus message number which matches the observations in eukaryotic cells (Fig. 4BC). However, in E. coli, the translation efficiency and message number are not strongly correlated (Fig. 4D). Why does this organism appear not to load balance? We demonstrate that the observed translation efficiency is consistent with the RLTO model, augmented by a ribosome-per-message limit. (See Sec. A.15.) Hausser et al. have proposed just such a limit, based on the ribosome footprint of mRNA molecules [10]. (See Sec. A.17.) Although this augmented model is consistent with central dogma regulation in E. coli, it is not a complete rationale. This proposed translation-rate limit could be circumvented by increasing the lifetime of E. coli messages which would increase the translation efficiency. Why the message lifetime is as short as observed will require a more detailed E. coli-specific analysis.
A.14 Discussion: Relation between load balancing and previous results